Temperature
(∘C)
Solubility
(g/100 g H2O)
20 11.0
30 23.0

You have prepared a saturated solution of X at 20∘C using 27.0 g of water. How much more solute can be dissolved if the temperature is increased to 30∘C?

Express your answer with the appropriate units.
Answer #1

Grams of water used for the saturated solution is 27.0 g. Now at 20 oC, the solubility of X is 11.0 g in 100 g of water., thus the amount of solute present in the 27.0 g water solution = 11*27/100 g = 2.97 g.

Now, if the temperature has raised to 30 oC, then the saturated solution will contain = 23*27/100 g = 6.21 g of solute in the solution at this temperature.

Thus the extra amount of solute that can be dissolved = (6.21-2.97) g = 3.24 g.